OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+2x+3=8 5x=8-3 5x=5 x=1 Check: 3(1+1)=-2(1-1)+6 6=6 true

OpenStudy (agreene):

3(x+1)=-2(x-1)+6 start by distributing 3 and -2 3x+3=-2x+2+6 combine like terms 3x+3=-2x+8 subtract 3 and add 2x to both sides 5x=5 divide by 5 on both sides x=1